# Learning Semantic Association Rules from Internet of Things Data

### v2

@@ 3.1 Input
$(i^{\prime}=(p^{\prime}\#z^{\prime}))$ (p1:).length $>$ 100 A node p1 has length bigger than 100
$(i^{\prime}=(m^{\prime}\#z^{\prime}))$ (s1:Sensor).value $<$ 10 A sensor s1 measures a value smaller than 10
@@ 3.2 Output
Let $I$ be a set of items. We define the following forms for an item, which are basic comparison operations: $\forall i^{\prime}\in I(((i^{\prime}=(p^{\prime}\#z^{\prime}))\lor(i^{\prime}=(m^{\prime}\#z^{\prime}))\lor(i^{\prime}=(v^{\prime}_{l}=l^{\prime}))\lor(i^{\prime}=(e^{\prime}_{l}=l^{\prime}))\lor(i^{\prime}=(v^{\prime}\rightarrow v^{\prime\prime}=e^{\prime})))$, with $p^{\prime}\in P$, $m^{\prime}\in M$, $v^{\prime},v^{\prime\prime}\in V$, $e^{\prime}\in E$, $l^{\prime},v^{\prime}_{l},e^{\prime}_{l}\in L$ where $v^{\prime}_{l}$ refers to a label mapped to a node with the ID $v^{\prime}$, and $e^{\prime}_{l}$ refers to a label mapped to an edge with the ID $e^{\prime}$. $z^{\prime}$ refers to a value that is either categorical or numerical, ##\## refers to one of the comparison operations with a truth value defined below:
#$\#_{categorical}(p,g)::=(p=g)|(p\neq g)|(p\in\{g\})|(p\notin\{g\})$$\#_{numerical}(p,g)::=(p=g)|(p\neq g)|(p>g)|(p<g)|(p\leq g)|(p\geq g)$
@@ 4.3 Rule Extraction from Autoencoders
Example: Figure 2 depicts an example rule extraction process. Assume that there are only two features in the input vector with 2 and 3 possible class values, namely $f_{1}=\{a,b\}$ and $f_{2}=\{c,d,e\}$. One-hot encoded version of $f_{1}$ and $f_{2}$ can be represented with 5 digits. Assume we want to test whether $f_{1}(a)$ implies a certain class value of $f_{2}$. Here we do a forward run on the trained Autoencoder with the input vector $[1,0,0.33,0.33,0.33]$ in which $f_{1}(a)$ is marked with 1, and equal probabilities are given for the values of $f_{2}$. We call this a $test\_vector$. Assume that the output is $[0.8,0.2,0.9,0.04,0,06]$. The third output digit that corresponds to $f_{2}(c)$ is bigger than the threshold, 0.8. Therefore, we say that $f_{1}(a)\rightarrow f_{2}(c)$.
Algorithm: The rule extraction algorithm is given in Algorithm 1. The parameters are the set of input vectors (input), a trained Autoencoder (ae), a similarity threshold (sim_threshold), and a maximum number of antecedents (antecedents) that the rules will contain. Based on the $antecedents$, in line 3, the algorithm creates combinations of features to be tested ($test\_ftrs)$, for instance, to test whether values of features $f_{1}$ and $f_{2}$ are associated with other features, a tuple of $(f_{1},f_{2})$ is created. Lines 4-13 go through each feature tuple ($ftr\_list$) in the test features and first create an initial test vector with all equal probabilities per feature (line 5). Line 6 marks feature values in the $ftr\_list$ with a probability of 1, and returns a list of test vectors ($test\_vectors$). Lines 7-13 perform a forward run per test vector and; i) check whether output probabilities for the marked features are higher than the given threshold (lines 9-10), ii) find features (other than marked features) that have higher probability than the given threshold, which are added to the rule list as consequences together with the marked features which are the antecedents (lines 11-13). The algorithm’s time complexity is $O(\binom{f}{a})$, where f represents the number of features and a denotes the maximum number of antecedents (see the section below for details).
@@
9:              if out_probs(ftr_list) $<$ sim_threshold then:
@@
12:                  if feature.prob() $>$ sim_threshold then:
@@
Line 2 initializes the $rules$ array, therefore it is $O(1)$.
@@
Line 4 iterates over the $test\_ftrs$. Therefore, the operations inside the loop are repeated $\binom{f}{a}$ times.
@@
Line 6 creates a set of vectors in which class values of the features in $ftr\_list$ are marked with 1. In the worst-case scenario, this step is linear over features when the $ftr\_list$ is equal to all of the features in the input dataset, hence, $O(f)$.
@@
Line 8 performs a forward pass with the given test vector. Since each forward pass performs a $softmax$ operation over the class values of features, this operation is linear over the number of features, $O(f)$, assuming that softmax is performed in $O(1)$.
Lines 9 and 10 perform a comparison operation to check whether probabilities inside the $out\_probs$ array that corresponds to the marked features are higher than a threshold or not. Assuming the worst-case scenario, this operation is repeated for each feature in the input data, $O(f)$.
@@ 5.1 Setup
  • Zhang’s Metric: This metric also considers the case in which the consequent appears alone in the transaction set, besides their co-occurrence, and therefore measures dissociation as well. A score of $>0$ indicates an association, 0 indicates independence and $<0$ indicates dissociation: $zm(X\rightarrow Y)=\frac{confidence(X\rightarrow Y)-confidence(X^{\prime}\rightarrow C)}{max(confidence(X\rightarrow Y),confidence(X^{\prime}\rightarrow Y))}$ in which $X^{\prime}$ refers to the absent of $X$ in the transaction set.
@@ 5.3 Experimental Results
Experiment 1.2: Execution Time. Figure 3 shows the effect of including semantics in the execution time of FP-Growth and Aerial (training + rule extraction time). The increase in the execution time of FP-Growth is 3-12 times while it is 2-3 times in Aerial and is more stable. However, since the semantic association rules have higher support and data coverage, a smaller number of them can have full data coverage (which is the case for Aerial and will be investigated in Experimental Setting 2). Therefore, we argue that the increment in the execution time is acceptable. Note that despite FP-Growth running faster with the parameters given in Table  4, it is strictly dependent on the preset minimum support threshold value and it runs slower for lower thresholds. This is investigated in Experiment 2.1.
@@ Appendix A Additional Experiments
FP-Growth (10) 43K |||| 472 0.23 |||| 0.22 0.24 |||| 0.23 0.96 |||| 0.96
FP-Growth (15) 130K |||| 7321 0.27 |||| 0.14 0.28 |||| 0.15 0.96 |||| 0.97
FP-Growth (20) 103K |||| 8974 0.41 |||| 0.19 0.43 |||| 0.2 0.95 |||| 0.97
Aerial (10) 123 |||| 109 0.31 |||| 0.24 0.33 |||| 0.27 0.94 |||| 0.91
Aerial (15) 547 |||| 940 0.31 |||| 0.27 0.37 |||| 0.31 0.88 |||| 0.89
Aerial (20) 554 |||| 2521 0.54 |||| 0.25 0.59 |||| 0.3 0.91 |||| 0.87
@@
FP-Growth (10) 11489 |||| 578 0.58 |||| 0.35 0.62 |||| 0.37 0.94 |||| 0.95
FP-Growth (15) 19447 |||| 2055 0.76 |||| 0.33 0.8 |||| 0.35 0.95 |||| 0.95
FP-Growth (20) 25421 |||| 5047 0.86 |||| 0.36 0.9 |||| 0.38 0.96 |||| 0.96
Aerial (10) 72 |||| 381 0.61 |||| 0.34 0.67 |||| 0.39 0.92 |||| 0.88
Aerial (15) 264 |||| 1300 0.54 |||| 0.35 0.6 |||| 0.42 0.9 |||| 0.87
Aerial (20) 1005 |||| 2551 0.59 |||| 0.39 0.65 |||| 0.45 0.91 |||| 0.88
@@
FP-Growth (10) 25 |||| 764 0.94 |||| 0.24 0.94 |||| 0.24 1 |||| 0.99
FP-Growth (15) 280 |||| 181 0.75 |||| 0.35 0.75 |||| 0.35 1 |||| 0.99
FP-Growth (20) 7220 |||| 2883 0.84 |||| 0.73 0.85 |||| 0.75 0.98 |||| 0.99
Aerial (10) 422 |||| 14 0.73 |||| 0.28 0.73 |||| 0.29 1 |||| 0.97
Aerial (15) 832 |||| 61 0.78 |||| 0.42 0.78 |||| 0.43 1 |||| 0.99
Aerial (20) 73 |||| 258 0.74 |||| 0.65 0.74 |||| 0.66 1 |||| 0.99

### v3

@@ 3.1 Input
$(i^{\prime}=(p^{\prime}\#z^{\prime}))$ (p1:).length >> 100 A node p1 has length bigger than 100
$(i^{\prime}=(m^{\prime}\#z^{\prime}))$ (s1:Sensor).value << 10 A sensor s1 measures a value smaller than 10
@@ 3.2 Output
Let II be a set of items. We define the following forms for an item, which are basic comparison operations: $\forall i^{\prime}\in I(((i^{\prime}=(p^{\prime}\#z^{\prime}))\lor(i^{\prime}=(m^{\prime}\#z^{\prime}))\lor(i^{\prime}=(v^{\prime}_{l}=l^{\prime}))\lor(i^{\prime}=(e^{\prime}_{l}=l^{\prime}))\lor(i^{\prime}=(v^{\prime}\rightarrow v^{\prime\prime}=e^{\prime})))$, with $p^{\prime}\in P$, $m^{\prime}\in M$, $v^{\prime},v^{\prime\prime}\in V$, $e^{\prime}\in E$, $l^{\prime},v^{\prime}_{l},e^{\prime}_{l}\in L$ where $v^{\prime}_{l}$ refers to a label mapped to a node with the ID $v^{\prime}$, and $e^{\prime}_{l}$ refers to a label mapped to an edge with the ID $e^{\prime}$. $z^{\prime}$ refers to a value that is either categorical or numerical, #\# refers to one of the comparison operations with a truth value defined below:
$_{categorical}(p,g)::=(p=g)|(p\neq g)|(p\in\{g\})|(p\notin\{g\})$ $_{numerical}(p,g)::=(p=g)|(p\neq g)|(p>g)|(p<g)|(p\leq g)|(p\geq g)$
@@ 4.3 Rule Extraction from Autoencoders
Example: Figure 2 depicts an example rule extraction process. Assume that there are only two features in the input vector with 2 and 3 possible class values, namely $f_{1}=\{a,b\}$ and $f_{2}=\{c,d,e\}$. One-hot encoded version of $f_{1}$ and $f_{2}$ can be represented with 5 digits. Assume we want to test whether $f_{1}(a)$ implies a certain class value of $f_{2}$. Here we do a forward run on the trained Autoencoder with the input vector $[1,0,0.33,0.33,0.33]$ in which $f_{1}(a)$ is marked with 1, and equal probabilities are given for the values of $f_{2}$. We call this a test_vectortest\_vector. Assume that the output is $[0.8,0.2,0.9,0.04,0,06]$. The third output digit that corresponds to $f_{2}(c)$ is bigger than the threshold, 0.8. Therefore, we say that $f_{1}(a)\rightarrow f_{2}(c)$.
Algorithm: The rule extraction algorithm is given in Algorithm 1. The parameters are the set of input vectors (input), a trained Autoencoder (ae), a similarity threshold (sim_threshold), and a maximum number of antecedents (antecedents) that the rules will contain. Based on the antecedentsantecedents, in line 3, the algorithm creates combinations of features to be tested (test_ftrs)test\_ftrs), for instance, to test whether values of features $f_{1}$ and $f_{2}$ are associated with other features, a tuple of $(f_{1},f_{2})$ is created. Lines 4-13 go through each feature tuple (ftr_listftr\_list) in the test features and first create an initial test vector with all equal probabilities per feature (line 5). Line 6 marks feature values in the ftr_listftr\_list with a probability of 1, and returns a list of test vectors (test_vectorstest\_vectors). Lines 7-13 perform a forward run per test vector and; i) check whether output probabilities for the marked features are higher than the given threshold (lines 9-10), ii) find features (other than marked features) that have higher probability than the given threshold, which are added to the rule list as consequences together with the marked features which are the antecedents (lines 11-13). The algorithm’s time complexity is $O(\binom{f}{a})$, where f represents the number of features and a denotes the maximum number of antecedents (see the section below for details).
@@
9:     if out_probs(ftr_list) << sim_threshold then:
@@
12:      if feature.prob() >> sim_threshold then:
@@
Line 2 initializes the rulesrules array, therefore it is $O(1)$.
@@
Line 4 iterates over the test_ftrstest\_ftrs. Therefore, the operations inside the loop are repeated $\binom{f}{a}$ times.
@@
Line 6 creates a set of vectors in which class values of the features in ftr_listftr\_list are marked with 1. In the worst-case scenario, this step is linear over features when the ftr_listftr\_list is equal to all of the features in the input dataset, hence, $O(f)$.
@@
Line 8 performs a forward pass with the given test vector. Since each forward pass performs a softmaxsoftmax operation over the class values of features, this operation is linear over the number of features, $O(f)$, assuming that softmax is performed in $O(1)$.
Lines 9 and 10 perform a comparison operation to check whether probabilities inside the out_probsout\_probs array that corresponds to the marked features are higher than a threshold or not. Assuming the worst-case scenario, this operation is repeated for each feature in the input data, $O(f)$.
@@ 5.1 Setup
  • Zhang’s Metric: This metric also considers the case in which the consequent appears alone in the transaction set, besides their co-occurrence, and therefore measures dissociation as well. A score of >0>0 indicates an association, 0 indicates independence and <0<0 indicates dissociation: $zm(X\rightarrow Y)=\frac{confidence(X\rightarrow Y)-confidence(X^{\prime}\rightarrow C)}{max(confidence(X\rightarrow Y),confidence(X^{\prime}\rightarrow Y))}$ in which $X^{\prime}$ refers to the absent of $X$ in the transaction set.
@@ 5.3 Experimental Results
Experiment 1.2: Execution Time. Figure 3 shows the effect of including semantics in the execution time of FP-Growth and Aerial (training + rule extraction time). The increase in the execution time of FP-Growth is 3-12 times while it is 2-3 times in Aerial and is more stable. However, since the semantic association rules have higher support and data coverage, a smaller number of them can have full data coverage (which is the case for Aerial and will be investigated in Experimental Setting 2). Therefore, we argue that the increment in the execution time is acceptable. Note that despite FP-Growth running faster with the parameters given in Table 4, it is strictly dependent on the preset minimum support threshold value and it runs slower for lower thresholds. This is investigated in Experiment 2.1.
@@ Appendix A Additional Experiments
FP-Growth (10) 43K $|$ 472 0.23 $|$ 0.22 0.24 $|$ 0.23 0.96 $|$ 0.96
FP-Growth (15) 130K $|$ 7321 0.27 $|$ 0.14 0.28 $|$ 0.15 0.96 $|$ 0.97
FP-Growth (20) 103K $|$ 8974 0.41 $|$ 0.19 0.43 $|$ 0.2 0.95 $|$ 0.97
Aerial (10) 123 $|$ 109 0.31 $|$ 0.24 0.33 $|$ 0.27 0.94 $|$ 0.91
Aerial (15) 547 $|$ 940 0.31 $|$ 0.27 0.37 $|$ 0.31 0.88 $|$ 0.89
Aerial (20) 554 $|$ 2521 0.54 $|$ 0.25 0.59 $|$ 0.3 0.91 $|$ 0.87
@@
FP-Growth (10) 11489 $|$ 578 0.58 $|$ 0.35 0.62 $|$ 0.37 0.94 $|$ 0.95
FP-Growth (15) 19447 $|$ 2055 0.76 $|$ 0.33 0.8 $|$ 0.35 0.95 $|$ 0.95
FP-Growth (20) 25421 $|$ 5047 0.86 $|$ 0.36 0.9 $|$ 0.38 0.96 $|$ 0.96
Aerial (10) 72 $|$ 381 0.61 $|$ 0.34 0.67 $|$ 0.39 0.92 $|$ 0.88
Aerial (15) 264 $|$ 1300 0.54 $|$ 0.35 0.6 $|$ 0.42 0.9 $|$ 0.87
Aerial (20) 1005 $|$ 2551 0.59 $|$ 0.39 0.65 $|$ 0.45 0.91 $|$ 0.88
@@
FP-Growth (10) 25 $|$ 764 0.94 $|$ 0.24 0.94 $|$ 0.24 1 $|$ 0.99
FP-Growth (15) 280 $|$ 181 0.75 $|$ 0.35 0.75 $|$ 0.35 1 $|$ 0.99
FP-Growth (20) 7220 $|$ 2883 0.84 $|$ 0.73 0.85 $|$ 0.75 0.98 $|$ 0.99
Aerial (10) 422 $|$ 14 0.73 $|$ 0.28 0.73 $|$ 0.29 1 $|$ 0.97
Aerial (15) 832 $|$ 61 0.78 $|$ 0.42 0.78 $|$ 0.43 1 $|$ 0.99
Aerial (20) 73 $|$ 258 0.74 $|$ 0.65 0.74 $|$ 0.66 1 $|$ 0.99